Introduction INGRID

has to confirm the stability of the neutrino beam flux and direction. For this purpose, we counted the number of neutrino events at each module from which the beam neutrino rate is monitored and the profile is reconstructed. Fig.?? shows the typical neutrino event at IN-GRID. Such a neutrino event in which there is at least one long track is counted.

[htbp]

[width=0.8]pdf/NuEvt00.pdf The typical neutrino event

INGRID adopt Multi Pixel Pho-